The perimeter of a rectangle is 46 ft. The length of the rectangle is 5 ft less than 3 times the width. Find the length of the rectangle.
Answer by checkley75(3666)

L=3W-5
2L+2W=46
2(3W-5)+2W=46
6W-10+2W=46
8W=46+10
8W=56
W=56/8
W=7 ANSWER FOR THE WIDTH.
L=3*7-5
L=21-5
L=16 ANSWER FOR THE LENGTH.
PROOF
2*16+2*7=46
32+14=46
46=46